# Advancing high-fidelity 3D and Texture Generation with 2.5D latents

### v1

@@ 4.4. Comparison with State-of-the-Art Methods
Table 1. Quantitative comparison for image-to-3D generation. We compare our method with existing approaches across geometric (CD, FS) and perceptual (PSNR, SSIM, LPIPS) metrics. Our method achieves strong performance, while using significantly less training data (26K). Notably, CraftsMan only reports geometric results (CD and FS) and does not model appearance, hence missing values in PSNR, SSIM, and LPIPS.
Figure 5. Qualitative comparison on image-to-3D generation. Please zoom in for detail.
Text-to-3D generation. As shown in Table 2, our method achieves the best overall performance across all three metrics: CLIP score, quality, and aesthetic scores. In particular, we outperform other approaches by a clear margin in terms of text-image alignment and perceptual quality. These trends are also evident in the qualitative results shown in Figure 6. Compared with Trellis-text-large and Kiss3DGen (Lin et al., 2025), our method generates 3D assets that more faithfully capture both the semantic content and fine-grained visual attributes described in the prompts. For example, in the “baby Groot” example, our model produces a character with more pronounced baby-like features—such as rounder eyes, a softer expression, and a more stylized, animated appearance—better matching the intended description. We attribute this to our unified 2.5D representation and effective adaptation of 2D diffusion priors, which together enable precise control over both appearance and structure. These results demonstrate the effectiveness of our approach in generating semantically grounded and visually appealing 3D assets from diverse natural language inputs.
@@
Table 2. Quantitative comparison of text-to-3D generation results across CLIP score, perceptual quality, and aesthetic rating. Our method outperforms all baselines on all three metrics.
@@
Figure 6. Qualitative comparison on text-to-3D generation. Please zoom in for detail.
Figure 7. Qualitative comparison of image and geometry conditioned texture generation. Our model perform the best in terms of texture fidelity. Please zoom in for detail.
Figure 8. Qualitative comparison of text and geometry conditioned texture generation. Compare to the baseline methods, our model better interprets the text condition and generates texture coherent with the given geometry conditions.
@@
Table 3. Qualitative comparison of text-to-texture generation results.
@@
Table 4. Qualitative comparison of image-to-texture generation results.
4.5. Ablation study
@@
Figure 9. Qualitative comparison of image-to-3D generation results (decoded from 2.5D latent to 3D Gaussian Splats) of our models trained with different strategy. The base model trained with Mixture-of-LoRA and image-text joint condition perform significantly better in producing results with better fidelity and 3D coherence. Please zoom in for detail.
@@ 6. Conclusion
Figure 10. Examples of image-and-geometry-to-texture generation with our framework. Please zoom in for detail.
Figure 11. Examples of image-to-3D generation with our framework. (a) We apply an extra refinement process to improve the surface and texture detail on the mesh decoder output using ISOMER (Wu et al., 2024). (b) The 3DGS results are rendered from decoder output directly. Please zoom in for detail.
@@ A.1. Mixture-of-LoRA
A.2. Extended Rotary Positional Embedding
@@
A.3. 3D Sparse Decoder for 3DGS and Mesh
@@
Mesh Decoder. The mesh decoder shares the same architectural backbone but uses 12 heads and a wider hidden size of 768. It predicts signed distance values for voxel-centered cubes following the FlexiCubes (Shen et al., 2023) representation. Two convolutional upsampling layers are applied post-decoding to increase spatial resolution and extract 0-level isosurfaces as meshes.
@@ Appendix B Inference Details
Figure 12. Qualitative comparison between the decoded 3DGS from generated 2.5D images (a) with or (b) without coordinate correction. The result with coordinate correction has better structural alignment with the generated multiview 2.5D images. Please zoom in for details.

### v2

@@ 4.4 Comparison with State-of-the-Art Methods
Table 1: Quantitative comparison for image-to-3D generation. We compare our method with existing approaches across geometric (CD, FS) and perceptual (PSNR, SSIM, LPIPS) metrics. Our method achieves strong performance, while using significantly less training data (26K). Notably, CraftsMan only reports geometric results (CD and FS) and does not model appearance, hence missing values in PSNR, SSIM, and LPIPS.
Figure 5: Qualitative comparison on image-to-3D generation. Please zoom in for detail.
Text-to-3D generation. As shown in Table 2, our method achieves the best overall performance across all three metrics: CLIP score, quality, and aesthetic scores. In particular, we outperform other approaches by a clear margin in terms of text-image alignment and perceptual quality. These trends are also evident in the qualitative results shown in Figure 6. Compared with Trellis-text-large and Kiss3DGen [21], our method generates 3D assets that more faithfully capture both the semantic content and fine-grained visual attributes described in the prompts. For example, in the “baby Groot” example, our model produces a character with more pronounced baby-like features—such as rounder eyes, a softer expression, and a more stylized, animated appearance—better matching the intended description. We attribute this to our unified 2.5D representation and effective adaptation of 2D diffusion priors, which together enable precise control over both appearance and structure. These results demonstrate the effectiveness of our approach in generating semantically grounded and visually appealing 3D assets from diverse natural language inputs.
@@
Table 2: Quantitative comparison of text-to-3D generation results across CLIP score, perceptual quality, and aesthetic rating. Our method outperforms all baselines on all three metrics.
@@
Figure 6: Qualitative comparison on text-to-3D generation. Please zoom in for detail.
Figure 7: Qualitative comparison of image and geometry conditioned texture generation. Our model perform the best in terms of texture fidelity. Please zoom in for detail.
Figure 8: Qualitative comparison of text and geometry conditioned texture generation. Compare to the baseline methods, our model better interprets the text condition and generates texture coherent with the given geometry conditions.
@@
Table 3: Qualitative comparison of text-to-texture generation results.
@@
Table 4: Qualitative comparison of image-to-texture generation results.
4.5 Ablation study
@@
Figure 9: Qualitative comparison of image-to-3D generation results (decoded from 2.5D latent to 3D Gaussian Splats) of our models trained with different strategy. The base model trained with Mixture-of-LoRA and image-text joint condition perform significantly better in producing results with better fidelity and 3D coherence. Please zoom in for detail.
@@ A.1 Mixture-of-LoRA
A.2 Extended Rotary Positional Embedding
@@
A.3 3D Sparse Decoder for 3DGS and Mesh
@@
Mesh Decoder. The mesh decoder shares the same architectural backbone but uses 12 heads and a wider hidden size of 768. It predicts signed distance values for voxel-centered cubes following the FlexiCubes [33] representation. Two convolutional upsampling layers are applied post-decoding to increase spatial resolution and extract 0-level isosurfaces as meshes.
@@ Appendix B Inference Details
Figure 12: Qualitative comparison between the decoded 3DGS from generated 2.5D images (a) with or (b) without coordinate correction. The result with coordinate correction has better structural alignment with the generated multiview 2.5D images. Please zoom in for details.